OpenStudy (anonymous):

The area of triangle ABC is 31 1/4 square centimeters. What is the measure of side b?

OpenStudy (anonymous):



OpenStudy (sachintha):

 First we need to find h.

OpenStudy (sachintha):

\[Area\ of\ a\ triangle=\frac{1}{2}*base *height\]

OpenStudy (sachintha):

We already know the base and area. So calculate h.

OpenStudy (anonymous):

How do I find the height? @sachintha

OpenStudy (sachintha):

\[31.25 cm^2 = \frac{1}{2}*12.5*h\]

OpenStudy (sachintha):

Can you calculate now?

OpenStudy (sachintha):

Give it a try?

OpenStudy (sachintha):

\[h=\frac{31.25*2}{12.5}\] \[h=5\]

OpenStudy (sachintha):

You can find b using trigonometry. \[sin\ 45^0=\frac{h}{b}\]

OpenStudy (sachintha):

\[sin\ 45^0=\frac{1}{\sqrt2}\]